A person rides up a lift to a mountaintop, but the person’s ears fail to “pop”—that is, the pressure of the inner ear does not equalize with the outside atmosphere.

The radius of each eardrum is 0.40 cm. The pressure of the atmosphere drops from 1.010 × 105 Pa at the bottom of the lift to 0.998 × 105 Pa at the top.

 

a. What is the pressure difference between the inner and outer ear at the top of the mountain?

b. What is the magnitude of the net force on each eardrum?


1
Expert's answer
2021-06-02T09:36:45-0400

a)


"\Delta P=(1.010-0.998)\cdot10^5=1200\ Pa"

b)


"F=\pi (0.004)^212000=0.60\ N"